# MOSPA: Human Motion Generation Driven by Spatial Audio

### v1

@@ 5.2 Ablation Study
Latent Dimension. The default latent dimension of MOSPA’s encoder transformer is 512. In our study, reducing it to 256 slightly increases the APD dou2023c value but also degrades the R-precision and the FID, leading to an overall decline in model performance, as illustrated in row 1 and row 2 in Tab. 3.
@@
Number of Diffusion Steps. We evaluated MOSPA with varying diffusion step numbers, reducing it from 1000 to 100 and further to 4, as detailed in rows 1, 4, and 5 of Tab. 3. Fewer steps slightly degrade the performance as shown by the increase in FID and degradation in diversity, thereby lower the upper limit of the power of the model.
@@ 6 Conclusion
In this paper, we introduce a task for enabling virtual humans to respond realistically to spatial auditory stimuli. While prior work maps audio to motion, it often neglects the spatial characteristics of spatial audio. We introduce a comprehensive SAM dataset, capturing human movement in response to spatial audio, and propose MOSPA, a diffusion-based generative model with an attention-based fusion mechanism. Once trained, MOSPA synthesizes diverse, high-quality motions that adapt to varying spatial audio input with binaural recording. Extensive evaluations show MOSPA achieves state-of-the-art performance on this task. Limitations and Future Works. 1) Physical Correctness: While MOSPA generates diverse and semantically plausible motions, it lacks physical constraints, which may lead to physically implausible artifacts. Integrating physics-based control methods dou2023c ; luo2023perpetual ; tessler2024maskedmimic ; huang2025modskill ; zhang2024physpt ; zhang2024incorporating ; pan2025tokenhsi ; wang2024sims could improve motion realism and embodiment fidelity (see Fig. 9 for spatial audio-driven humanoid robot control). Body Modeling: This work focuses on body motion and omits finer-grained components such as hand gestures and facial expressions supported by SMPL-X pavlakos2019expressive . Extending the model to full-body motion generation lu2023humantomato ; lin2023motion ; xu2025intermimic ; pi2025coda ; bian2025motioncraft —including hand motions—remains an important direction for future research. Scene Awareness: The current framework does not incorporate awareness of surrounding environments or physical scene geometry, limiting its ability to produce scene-consistent or contact-aware motions. Future extensions could integrate scene representations or affordance prediction cong2024laserhuman ; wang2021scene ; wang2024move ; cen2024generating ; wang2022towards with spatial audio signals to enhance human motion generation.
@@ Acknowledgements
We are grateful to Yiduo Hao, Chuan Guo, Chen Wang and Peter Qingxuan Wu for their insightful discussions and constructive feedback, which have been helpful to the development of this research.
@@ Appendix A Dataset Statistics
Three general motion genres are covered in the SAM dataset—dull, neutral, and sensitive. The intensity and reaction speed decreased from sensitive to neutral to dull. The motion types listed in Tab. A5 can thus be further divided into motion types with genres, as listed in Tab. A6, A7, A8.
@@
look at nod head prey
@@ Appendix B Implementation Details
B.1 Audio Features
We present more details regarding the audio features extracted. The full list of audio features and their corresponding dimensions are listed in Tab. B9. The final shape of the audio features $\mathbf{a}$ is thus $\mathbf{a}\in\mathbb{R}^{T\times 2272}$, where T=240 is the number of frames in each motion sequence.

### v2

@@ 5.2 Ablation Study
Latent Dimension. The default latent dimension of MOSPA’s encoder transformer is 512. In our study, reducing it to 256 slightly increases the APD dou2023c value but also degrades the R-precision and the FID, leading to an overall decline in model performance, as seen in row 1 and row 2 in Tab. 3.
@@
Number of Diffusion Steps. We evaluated MOSPA with varying diffusion step numbers, reducing it from 1000 to 100 and further to 4, as detailed in rows 1, 4, and 5 of Tab. 3. Fewer steps slightly degrade the performance as shown by the increase in FID and degradation in diversity, thereby lowering the upper limit of the power of the model.
@@
Table 4: Ablation study on the effect of MFCC davis1980comparison and tempogram grosche2010cyclic features.
MFCC Tempogram $\text{R-precision}\uparrow$ $\text{FID}\downarrow$
Top-1 Top-2 Top-3
$\mathbf{0.937^{\pm 0.005}}$ $\mathbf{0.984^{\pm 0.002}}$ $\mathbf{0.996^{\pm 0.001}}$ $\mathbf{7.981}$
$\times$ $0.907^{\pm 0.004}$ $0.967^{\pm 0.002}$ $0.983^{\pm 0.002}$ $9.070$
$\times$ $0.917^{\pm 0.004}$ $0.982^{\pm 0.002}$ $0.994^{\pm 0.001}$ $10.786$
We evaluate the contribution of the extracted audio features by conducting an ablation study on the effectiveness of MFCC davis1980comparison and tempogram grosche2010cyclic features. As shown in Tab. 4, improvements in FID and R-precision—two key metrics for assessing generative quality and correspondence—demonstrate their significance in model.
@@ 6 Conclusion
This introduces a novel task for enabling virtual humans to respond realistically to spatial auditory stimuli. We present a comprehensive SAM dataset, capturing human movement in response to spatial audio, and propose MOSPA, a diffusion-based generative model with an attention-based fusion mechanism. Once trained, MOSPA synthesizes diverse, high-quality motions that adapt to varying spatial audio inputs with binaural recording. Extensive evaluations show MOSPA achieves state-of-the-art performance on this task. Limitations and Future Works. Physical Correctness: While MOSPA generates diverse and semantically plausible motions, it lacks physical constraints, which may lead to physically implausible artifacts. Integrating physics-based control methods dou2023c ; luo2023perpetual ; tessler2024maskedmimic ; huang2025modskill ; zhang2024physpt ; zhang2024incorporating ; pan2025tokenhsi ; wang2024sims could improve motion realism and embodiment fidelity (see Fig. 9 for spatial audio-driven humanoid robot control). Body Modeling: This work focuses on body motion and omits finer-grained components such as hand gestures and facial expressions supported by SMPL-X pavlakos2019expressive . Extending the model to full-body motion generation lu2023humantomato ; lin2023motion ; xu2025intermimic ; pi2025coda ; bian2025motioncraft —including hand motions—remains an important direction for future research. Scene Awareness: The current framework does not incorporate awareness of surrounding environments or physical scene geometry, limiting its ability to produce scene-consistent or contact-aware motions. Future extensions could integrate scene representations or affordance prediction cong2024laserhuman ; wang2021scene ; wang2024move ; cen2024generating ; wang2022towards with spatial audio signals to enhance human motion generation.
@@ Acknowledgements
This work is partly supported by the Innovation and Technology Commission of the HKSAR Government under the ITSP-Platform grant (Ref: ITS/335/23FP) and the InnoHK initiative (TransGP project). Part of the research was conducted in the JC STEM Lab of Robotics for Soft Materials, funded by The Hong Kong Jockey Club Charities Trust. We are grateful to Yiduo Hao, Chuan Guo, Chen Wang, Zitong Lan, and Peter Qingxuan Wu for their insightful discussions and constructive feedback, which have been helpful to the development of this research.
@@ Appendix A Dataset Statistics
Three general motion genres are covered in the SAM dataset—dull, neutral, and sensitive. The intensity and reaction speed decreased from sensitive to neutral to dull. The motion types listed in Tab. A6 can thus be further divided into motion types with genres, as listed in Tab. A7, A8, A9.
@@
look at nod head pray
@@ Appendix B Implementation Details
B.1 More Implementation Details
We present more details regarding the audio features extracted. The full list of audio features and their corresponding dimensions are listed in Tab. B10. During feature extraction of the audio, we use a sampling rate of 30,720 Hz, an FFT window length of 2048, and a hop length of 256. The Short-Time Fourier Transform (STFT) windowing function is the Hann window blackman1958measurement . The final shape of the audio features $\mathbf{a}$ is thus $\mathbf{a}\in\mathbb{R}^{T\times 2272}$, where T=240 is the number of frames in each motion sequence.